40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 960     Entry time: Wed Sep 17 19:13:47 2008
Author: Alberto 
Type: Update 
Category: General 
Subject: ABSL: status 
I installed the setup for measuring TEM01/10 on the X end table.
I'm leaving. I shut the laser, flipped down the flipper mirror, disconnected the pzt drive signal from the laser.
For Jenne. The power cable for the Guralps' board is now connected to the PDH box on my instruments cart but you can take it.
ELOG V3.1.3-